ADVANCED DYNAMICS — PHY 4241/5227 SOLUTIONS – SET 8

(25) One finds

( \begin{matrix}x^{′ 1} \\ i x^{′ 0} \end{matrix} ) = ( \begin{matrix}cosh(ζ) & i sinh(ζ) \\ −i sinh(ζ) & cosh(ζ) \end{matrix} ) ( \begin{matrix}x^{1} \\ i x^{0} \end{matrix} )

and in components

x^{′ 1} = cosh(ζ) x^{1} − sinh(ζ) x^{0}
i x^{′ 0} = −i sinh(ζ) x^{1} + i cosh(ζ) x^{0}

or, equivalently,

x^{′ 0} = + cosh(ζ) x^{0} − sinh(ζ) x^{1}
x^{′ 1} = − sinh(ζ) x^{0} + cosh(ζ) x^{1}

( \begin{matrix}x^{′ 1} \\ x^{′ 0} \end{matrix} ) = ( \begin{matrix}cosh(ζ) & − sinh(ζ) \\ − sinh(ζ) & cosh(ζ) \end{matrix} ) ( \begin{matrix}x^{1} \\ x^{0} \end{matrix} )